# NeuroPath: A Neural Pathway Transformer for Joining the Dots of Human Connectomes

### v1

@@ 1 Introduction
Figure 1: Planting a novel multivariate SC-FC coupling mechanism to explainable deep model. Orange box: The structural connectivity (SC) denoted by grey links represents the strength of neurological fiber that physically connects two brain regions. SC is relatively static given the neural activities are transient, e.g. cognitive tasks. Green box: The functional connectivity (FC) is commonly considered as the brain network topology [48] since SC is static for different cognitive tasks. The overlapping area of orange and green boxes: the multivariate SC-FC coupling mechanism, where a neural pathway (detour) is constructed by multiple SC links to support one FC link. Grey box: NeuroPath Transformer using a new MHSA module filtered by adjacency matrices has expressive power for multi-hop detours.
@@ 2.1 Motivations
Massive inter-subject variations of FC often predominate the intrinsic task-relevant patterns. Take the HCP data [9] as an example in which the data heterogeneity issue has been harmonized by using standardized imaging protocol. As shown in Fig. 2, we first identify a set of brain regions which exhibit significant difference between resting stage and VISMOTOR tasking [9] in a paired t-test using detour degree and FC degree [^1] Detailed steps of paired t-tests can be found in Appendix., respectively, where the identified regions are colored based on p-values. Next, we examine the location of these regions in each functional sub-networks (manually defined using neuroscience domain knowledge). For comparison, we display the overlap ratio (y-axis) between the number identified regions and the total number of regions in each sub-network (x-axis) at the bottom of Fig. 2. Current neuroscience findings suggest that resting state is relevant to default mode network (red circle) while VISMOTOR task is associated with visual (orange circle) and sensorimotor areas (green circle). In this regard, the detour degree achieves overall higher discrimination power than FC to the extent that most of identified regions are aligned with neuroscience finding with much less false positives in other non-relevant regions. The evidence shown in Fig. 2 underscores the importance of including SC detour in finding putative functional biomarkers, prompting us to further integrate the SC-FC coupling mechanism into the design of NeuroPath.
@@
2.2 Relevant Machine Learning Work for Brain
Connectomes
@@ 3 NeuroPath – A Biological-Inspired Transformer for Human Connectomes
Note that $\mathbf{D}^{h}$ avoids finding all simple paths by our model, it significantly reduces computational costs with sufficient expressive power for neural pathways as discussed in Sec. 3.2.
@@ 3.1 Network Architecture of NeuroPath
NeuroPath is designed with twin branch as shown in Figure 3, where two branches in the model have identical frameworks of the multi-head self-attention (MHSA) from the Transformer encoder [56]. This twin branch is designed for feature representation to fuse SC and FC information by being consistent between FC link representation and SC detour representation. Given a set of learnable parameters $\{\bar{\mathbf{W}},\hat{\mathbf{W}}\in\mathbb{R}^{(HC)\times C},\boldsymbol{\bar{\alpha}}_{1},\dots,\boldsymbol{\bar{\alpha}}_{H},\boldsymbol{\bar{\beta}}_{1},\dots,\boldsymbol{\bar{\beta}}_{H},\boldsymbol{\bar{\gamma}}_{1},\dots,\boldsymbol{\bar{\gamma}}_{H},$ $\boldsymbol{\hat{\alpha}}_{1},\dots,\boldsymbol{\hat{\alpha}}_{H},\boldsymbol{\hat{\beta}}_{1},\dots,\boldsymbol{\hat{\beta}}_{H},\boldsymbol{\hat{\gamma}}_{1},\dots,\boldsymbol{\hat{\gamma}}_{H}\in\mathbb{R}^{C\times C}\}$, where MHSA has $H$ heads and feature dimension $C$, the branch TD-MHSA is then defined as
@@
3.2 Expressive Power Analysis for NeuroPath
@@
By choosing proper weights, TD-MHSA at least has half of the expressive power of the path neural network (PathNN) [40] for $H$-hop neural pathways of the brain network.
Considering PathNN is graph expressive, Fact 3.1 indicates the power of our NeuroPath for high-order graph substructure. PathNN models all simple paths has $O(N^{H})$ computational complexity by sorting all paths in advance. In contrast, our NeuroPath can achieve half of the expressive power using only $O(N^{2})$ computational complexity and does not require any pre-computation of high-order topology.
Regarding the neuroscience knowledge that FC is relatively denser than SC with the evidence of significant indirect SC links [39], FC adjacency is also denser than the TD adjacency according to the Def. 3.1. Meanwhile, the consistency constraint $\mathcal{L}_{TD}$ cooperates with the twin branch so that FC-MHSA produces the expressive node feature representation as well. Consequently, existing states of FC-MHSA can hold the expressive power of TD-MHSA by a consistent feature representation, and hence benefit the prediction by this SC-FC coupling.
@@ 4.3 Zero-shot Learning
4.4 Ablation Studies
@@ A.1 Accessibility
All data can be accessible via internet (HCPA [^2] https://www.humanconnectome.org/, UKB [^3] https://www.ukbiobank.ac.uk/, ADNI [^4] https://adni.loni.usc.edu/, OASIS [^5] https://sites.wustl.edu/oasisbrains/). The licenses to obtain those data can also be accessed on the websites. The preprocessing algorithm is introduced in the next section. The codes and data split settings can be acquired via this anonymous code repository [^6] https://anonymous.4open.science/r/neuro_detour-47E1/.
@@ A.4 Detailed Steps of Paired t-test
A.5 Proof of Fact
Fact 3.1 By choosing proper weights, TD-MHSA at least has half of the expressive power of the path neural network (PathNN) [40] for $H$-hop neural pathways of the brain network.
@@ A.5 Proof of Fact
Taking Eq. 7 and Lemma A.1 together, $f_{TD}(\mathbf{X})_{i}$ weights all nodes of half of the neural pathways by sorting results $\mathbf{P}^{H}_{i}$. By assigning each node with the learnable weights $\mathbf{S}_{ij}\boldsymbol{\bar{\gamma}}_{h}\bar{\mathbf{W}}_{i\sim j}$ since $\mathbf{S}_{ij}$ is a scalar that can be moved aside to matrices, each path weights is thus represented by this weight consists of learnable parameters and can be extracted to explain the contribution by neural pathways to any downstream application.
With properly training the model, an expressive pathway representation can be achieved with half of the expressive power of PathNN for neural pathways, that is also aggregating nodes of neural pathways followed by a combination layer. This can finish the proof. ∎
A.6 Computational Costs of Models Tested in the Paper
Our experiments are run on a local platform that has dual Intel(R) Xeon(R) Gold 6448Y CPUs and four NVIDIA RTX 6000 Ada GPUs. All SOTA models in our experiments used the default hyperparameters except for Graphormer which is implemented by us. Detailed settings can be found in the released codes. The computational costs can be indicated by the learnable parameter amount as listed in Table 6.
Table 6: Model computational costs.
BrainGNN BNT BolT Graphormer NAGphormer NeuroPath
Param. number 7.30M 1.57M 1.58M 0.30M 0.26M 0.69M
A.7 Limitations and Discussions
The limitations of NeuroPath are two folds. First, the expressive power is limited on the half number of neural pathways in theory. This limitation is amplified in dynamic data that has a higher degree of connections leading to more pathways, where the performance is relatively lower than in static data as listed in the tables of our experiments. Although our detour adjacency matrix avoids finding all simple paths in terms of computational complexity, modeling a brain network with all neural pathways could make NeuroPath more fundamental. Second is the limited size of disease datasets used in the experiments.
A.8 Societal Impact

### v3

@@ 1 Introduction
Figure 1: Planting a novel multivariate SC-FC coupling mechanism to explainable deep model. Orange box: The structural connectivity (SC) denoted by grey links represents the strength of neurological fiber that physically connects two brain regions. SC is relatively static given the neural activities are transient, e.g. cognitive tasks. Green box: The functional connectivity (FC) is commonly considered as the brain network topology [48] since SC is static for different cognitive tasks. The overlapping area of orange and green boxes: the multivariate SC-FC coupling mechanism, where a neural pathway (detour) is constructed by multiple SC links to support one FC link. Grey box: NeuroPath Transformer using a new MHSA module filtered by adjacency matrices emits the representation of multi-hop detours.
@@ 2.1 Motivations
Massive inter-subject variations of FC often predominate the intrinsic task-relevant patterns. Take the HCP data [9] as an example in which the data heterogeneity issue has been harmonized by using standardized imaging protocol. As shown in Fig. 2, we first identify a set of brain regions which exhibit significant difference between resting stage and VISMOTOR tasking [9] in a paired t-test using detour degree and FC degree [^1] Detailed steps of paired t-tests can be found in Appendix., respectively, where the identified regions are colored based on p-values. Next, we examine the location of these regions in each functional sub-networks (manually defined using neuroscience domain knowledge). For comparison, we display the overlap ratio (y-axis) between the number identified regions and the total number of regions in each sub-network (x-axis) at the bottom of Fig. 2. Current neuroscience findings suggest that resting state is relevant to default mode network (red circle) while VISMOTOR task is associated with visual (orange circle) and sensorimotor areas (green circle). In this regard, the detour degree achieves overall higher discrimination power than FC to the extent that most of identified regions are aligned with neuroscience finding with much less false positives in other non-relevant regions. The same findings are observed on diverse populations by gender in Appendix Fig. 8. The evidence shown in Fig. 2 underscores the importance of including SC detour in finding putative functional biomarkers, prompting us to further integrate the SC-FC coupling mechanism into the design of NeuroPath.
@@
2.2 Relevant Machine Learning Work for
Brain Connectomes
@@ 3 NeuroPath – A Biological-Inspired Transformer for Human Connectomes
Note that $\mathbf{D}^{h}$ avoids finding all simple paths by our model, it significantly reduces computational costs with sufficient power for the neural pathways representation as discussed in Sec. 3.2.
@@ 3.1 Network Architecture of NeuroPath
NeuroPath is designed with twin branch as shown in Figure 3, where two branches in the model have identical frameworks of the multi-head self-attention (MHSA) from the Transformer encoder [56]. This twin branch is designed for feature representation to fuse SC and FC information by being consistent between FC link representation and SC detour representation. Given a set of learnable parameters $\bar{\mathbf{W}},\hat{\mathbf{W}}\in\mathbb{R}^{(HC)\times C}$ and $\boldsymbol{\bar{\alpha}}_{h},\boldsymbol{\bar{\beta}}_{h},\boldsymbol{\bar{\gamma}}_{h},\boldsymbol{\hat{\alpha}}_{h},\boldsymbol{\hat{\beta}}_{h},\boldsymbol{\hat{\gamma}}_{h}\in\mathbb{R}^{C\times C}$ where $h=1,\dots,H$, $H$ is head number of MHSA, and $C$ denotes feature dimension, the branch TD-MHSA is then defined as
@@
3.2 Neural Pathway Representation by NeuroPath
@@
The top pathway representations are obtained by $\arg\max_{j,h}\left(\frac{1}{h}\sum_{j\in\mathbf{p}}\mathbf{S}_{ij}\boldsymbol{\bar{\gamma}}_{h}\bar{\mathbf{W}}_{i\sim j}\right)$, where $\mathbf{S}$ denotes the softmax of self-attention, $\mathbf{p}\subset\mathbf{P}^{H}_{i}$ is a set of node index of a path and $\mathbf{P}^{H}_{i}$ is the node collection of neural pathways within $H$-hop starting at $i$-th node.
Fact 3.1 indicates the power of our NeuroPath for high-order graph substructure. PathNN [40] models all simple paths has $O(N^{H})$ computational complexity by sorting all paths in advance. In contrast, NeuroPath can model important paths using only $O(N^{2})$ computational complexity and does not require any pre-computation of high-order topology.
Regarding the neuroscience knowledge that FC is relatively denser than SC with the evidence of significant indirect SC links [39], FC adjacency is also denser than the TD adjacency according to the Def. 3.1. Meanwhile, the consistency constraint $\mathcal{L}_{TD}$ cooperates with the twin branch so that FC-MHSA produces the expressive node feature representation as well. Consequently, existing states of FC-MHSA can hold the same power of TD-MHSA by a consistent feature representation, and hence benefit the prediction by this SC-FC coupling.
@@ 4 Experimental Results
Table 3: Conclusion of performance: Average rank of methods on different scenarios and evaluation metrics. ‘Bold’ and ‘underline’ denote the first and the second place, respectively.
MLP GCN BrainGNN BNT BolT Graphormer NAGphormer NeuroPath
HCPA 4.0 4.5 7.0 2.8 2.5 8.0 5.3 2.0
UKB 3.0 3.75 7.0 5.0 2.3 8.0 5.3 1.8
ADNI 6.9 4.4 4.1 2.1 5.8 4.6 5.8 1.6
OASIS 3.3 5.3 4.4 2.8 6.5 6.4 5.0 2.5
@@ 4.3 Zero-shot Learning
4.4 Ablation of Hyperparameter $H$
4.5 Ablation of Model Architecture
@@
4.6 Ablation of Model Depth
Table 6: F1 scores by models with different layer numbers on four datasets. ‘Bold’ and ‘underline’ denote the first and the second place of the average rank, respectively.
HCPA Rank UKB Rank
Layer # 4 8 16 4 8 16
BNT 91.81 93.41 93.28 3.67 88.63 96.32 97.45 3.00
BolT 97.01 97.81 88.23 2.33 81.36 89.20 89.84 4.00
Graphormer 64.08 47.01 50.84 5.00 43.42 43.44 59.46 5.00
NAGphormer 96.89 97.26 97.22 2.33 99.24 98.95 99.20 2.00
NeuroPath 97.76 97.72 96.60 1.67 99.59 99.61 99.44 1.00
ADNI Rank OASIS Rank
BNT 76.39 75.91 77.28 3.67 85.32 85.96 85.21 3.33
BolT 75.93 78.67 78.23 2.67 85.30 84.55 85.55 3.67
Graphormer 78.58 74.12 74.12 4.00 84.45 83.87 83.87 5.00
NAGphormer 75.86 77.15 78.44 3.00 86.05 86.49 85.78 1.67
NeuroPath 78.93 78.42 78.32 1.67 86.16 86.77 85.78 1.00
Performance comparison between models with different layer numbers is listed in Table 6. Most of competing models dropped on performance when increasing the layer number to 16, e.g., Graphormer dropped under 50 on HCPA and UKB datasets. In contrast, NeuroPath is more stable than others and has the best average ranking.
4.7 Ablation of Graph Building
Performance comparison between models with different edge threshold to construct brain network is listed in Table 7. Similarly, existing methods dropped on performance when graph is too dense or too sparse as the threshold of Pearson coefficient changed, which is an empirical hyperparameter. NeuroPath can keep the best average ranking in this case.
@@ 4.8 Pattern of Neural Pathway Contributing to Prediction
Table 7: F1 scores by models with different FC thresholds to build brain networks. ‘Bold’ and ‘underline’ denote the first and the second place of the average rank, respectively.
HCPA Rank UKB Rank
FC threshold 0.3 0.5 0.7 0.3 0.5 0.7
BNT 95.73 92.57 84.51 4.00 76.41 98.64 94.46 4.33
BolT 87.02 95.78 94.68 3.00 86.98 99.29 87.04 3.67
Graphormer 90.41 53.05 88.43 4.33 97.76 86.54 96.73 3.67
NAGphormer 96.08 94.76 96.85 2.33 97.80 99.22 98.78 2.33
NeuroPath 97.57 95.09 97.32 1.33 99.27 99.59 99.15 1.00
ADNI Rank OASIS Rank
BNT 77.74 80.16 77.92 1.67 85.14 85.32 86.05 3.67
BolT 74.33 76.68 76.53 4.00 84.98 84.91 84.67 4.67
Graphormer 75.82 77.78 75.17 3.33 86.23 85.44 87.15 2.00
NAGphormer 72.55 75.40 77.29 4.67 86.32 83.87 85.78 3.67
NeuroPath 78.36 77.35 79.49 1.67 86.59 87.02 86.13 1.33
Table 8: Computational complexity in our experiments, where computing time is the average on UKB dataset with unit the millisecond per graph data.
Param # Preproc. Train Test
BrainGNN 7.30M - 7.24 2.61
BNT 1.57M - 1.82 0.64
BolT 1.58M - 3.83 1.83
Graphormer 0.30M 270 2.79 0.90
NAGphormer 0.26M 40 3.92 1.85
NeuroPath 0.69M - 1.61 0.67
@@
4.9 Computational Costs
The computational costs can be indicated by the practical running time and the learnable parameter amount as listed in Table 8. Although Graphormer and NAGphormer are two models with lower parameter numbers than NeuroPath, they have slower training and testing than NeuroPath with pre-processing leading to more computing time. This demonstrates the efficiency of NeuroPath.
@@ A.1 Accessibility
All data can be accessible via internet (HCPA [^2] https://www.humanconnectome.org/, UKB [^3] https://www.ukbiobank.ac.uk/, ADNI [^4] https://adni.loni.usc.edu/, OASIS [^5] https://sites.wustl.edu/oasisbrains/). The licenses to obtain those data can also be accessed on the websites. The preprocessing algorithm is introduced in the next section. The codes and data split settings can be acquired via this code repository [^6] https://github.com/chrisa142857/neuro_detour.
@@ A.3 Data Profiles
One major phenotype of brain networks is the age of the subject as shown in Fig. 7. Accordingly, our experiments are setup with a wide range of ages.
@@
Figure 7: The age distribution of HCPA and ADNI datasets, where blue bars are histograms of age, and orange curves denote the data percentage accumulation.
@@ A.4 Detailed Steps of Paired t-test
A.5 Paired t-test on diverse populations by gender on HCPA dataset
As shown in 8, we run the same $t$-test as in Fig.2 in the main text for the degree of our topological detour across diverse populations by gender. By comparing the detour degree with the FC degree, we can draw the same conclusion as in Sec 2.1. It is worth noting that the male group shows a lower detour degree, i.e., fewer detour pathways, than the female group in the HCPA dataset.
Figure 8: The same $t$-test as Fig.2 in the main text runs on two diverse populations by gender among HCPA dataset.
A.6 Proof of Fact
Fact 3.1 The top pathway representations are obtained by $\arg\max_{j,h}\left(\frac{1}{h}\sum_{j\in\mathbf{p}}\mathbf{S}_{ij}\boldsymbol{\bar{\gamma}}_{h}\bar{\mathbf{W}}_{i\sim j}\right)$, where $\mathbf{S}$ denotes the softmax of self-attention, $\mathbf{p}\subset\mathbf{P}^{H}_{i}$ is a set of node index of a path and $\mathbf{P}^{H}_{i}$ is the node collection of neural pathways within $H$-hop starting at $i$-th node.
@@ A.6 Proof of Fact
Taking Eq. 7 and Lemma A.1 together, $f_{TD}(\mathbf{X})_{i}$ weights all nodes of half of the neural pathways by sorting results $\mathbf{P}^{H}_{i}$. By assigning each node with the learnable weights $\mathbf{S}_{ij}\boldsymbol{\bar{\gamma}}_{h}\bar{\mathbf{W}}_{i\sim j}$ since $\mathbf{S}_{ij}$ is a scalar that can be moved aside to matrices, each path weights is thus represented by this weight consists of learnable parameters and can be extracted to explain the contribution by neural pathways to any downstream application. This can finish the proof. ∎
A.7 Hardwares
Our experiments are run on a local platform that has dual Intel(R) Xeon(R) Gold 6448Y CPUs and four NVIDIA RTX 6000 Ada GPUs. All SOTA models in our experiments used the default hyperparameters except for Graphormer which is implemented by us. Detailed settings can be found in the released codes.
A.8 Limitations and Discussions
The limitations of NeuroPath are two folds. First, the top path representation is limited on the half number of neural pathways in theory. This limitation is amplified in dynamic data that has a higher degree of connections leading to more pathways, where the performance is relatively lower than in static data as listed in the tables of our experiments. Although our detour adjacency matrix avoids finding all simple paths in terms of computational complexity, modeling a brain network with all neural pathways could make NeuroPath more fundamental. Second is the limited size of disease datasets used in the experiments.
A.9 Societal Impact
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In Abstract and Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Appendix
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section 3.2 and Appendix
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Appendix
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: In Appendix
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 4 and Appendix
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Section 4
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: [N/A]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Appendix
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In Section 4
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.